how to solve this? |2x+3|=3
Answer by Boreal(15235)

For absolute value problems
Set the equation without brackets = 3, as if they weren't there.
Then set the negative of it = 3. The negative in absolute value brackets will be 3 once the brackets are removed.
2x+3=3
x=0, 1 solution.
-2x-3=3
-2x=6
x=-3
Check solutions
|3|=3
|-6+3|=|-3| =3